OpenStudy (anonymous):

Find the angles in a triangle if the measures of the angles are a.) x, 2x, and 3x b.) x-10,x, and x+10 I'm in grade nine. Exam on monday! I have no idea what this question is talking about, oh yea, is there a way to change my username? I made it a year ago and I want to change it to TrueDirectioner or VasHappening?

OpenStudy (anonymous):

Thanks Mertsj!

OpenStudy (mertsj):

Do you know that when you add the three angles of a triangle you always get 180?

OpenStudy (anonymous):

yup. I'm not that dumb lol

OpenStudy (mertsj):

So they said the three angles of the triangle are x, 2x and 3x. So you should write and solve the equation that says the sum of those 3 is 180

OpenStudy (anonymous):

x+2x+3x=180 6x=180 *180 divided by 6* x= 30

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

Yes. x = 30 and 2x = ? and 3x= ?

OpenStudy (anonymous):

wait is that right?

OpenStudy (anonymous):

2x=60

OpenStudy (anonymous):

3x=90

OpenStudy (mertsj):

Do those 3 add up to 180?

OpenStudy (anonymous):

30 +60+90 = 180!

OpenStudy (mertsj):

So now you have found the 3 angles of that triangle. And the b example is similar.

OpenStudy (mertsj):

I'm not in a grade.